lacques and Kyoko are partners at a management consulting firm. They are trying to determine which of them has a comparative advantage in creating the 25 slides required for a sales pitch to a prospective dient lacques can create 20 slides per hour. For other activties, he can bill clients $400 per hour. Jacquess opportunity cost of creating slides is Per side. Kyokos opportunity ost of creating sides is 35% higher than Jacquess. However, as the unior partner, her billing rate is 30% lower. Based on all of these facts has a comparative advantage in creating slides Grade tt Now Save& continue Continue wthout saving
Answer #1

Opportunity cost of creating slides is 400/20=$20

kyoko opportunity cost of slide=1.35*20=27$ and billing rate=0.7(400)=280

thus opportunity cost of Kyoko is 280/27=10.37

jacques has lower opportunity cost in creating slides thus Jacques has comparative advantage in slides